OpenStudy (anonymous):

a car is travelling at 30 m/s. suddenly the driver sees a danger ahead and decides to do an emergency stop. the driver's reaction time is 0.6 s. calculate how far the car travels before the driver starts pressing the brake pedal

OpenStudy (anonymous):

30 x 0.6 = 18 m

OpenStudy (anonymous):

t = 0.6s v = 30 m/s v = s / t, hence s = v . t so s = 0.6 x 30 = 18 metre